OpenStudy (anonymous):

Could I have some help with this Algebra 2 question? I am not for sure how to do this and every number I come up with makes no sense. 3 (^5sqrt x+2^3)+3 =27

OpenStudy (amistre64):

can you show how you are getting your numbers?

OpenStudy (anonymous):

I subtracted by three then I divided 24 by 3 and got 8

OpenStudy (amistre64):

^5sqrt x+2^3, this is not very readable perhaps you mean: 5rt(x+2^3) ? or is it: 5rt(x) + 2^3 ?

OpenStudy (anonymous):

after that I got confused what happened next

OpenStudy (amistre64):

2rt = sqrt 3rt = cbrt 4rt 5rt etc ... these are easier to read then trying the ^n sqrt

OpenStudy (anonymous):

the first one you typed was right! I couldn't figure out how to type it.

OpenStudy (amistre64):

3k+3 =27 3k =24 k = 8 then unroot it k^(1/5) = 8^(1/5) assuming k is a 5rt radical

OpenStudy (amistre64):

or if k = 5rt(x) + 2^3 then we continue by subtracting 8 and then worry about the rt

OpenStudy (anonymous):

\[3\sqrt[5]{x+2^{3}}+3=27\]

OpenStudy (anonymous):

I get stuck at 5rt(x+2)^3 = 8

OpenStudy (amistre64):

which is it: \[a)~~~\sqrt[5]{x+(2)^3}\]or \[b)~~~\sqrt[5]{(x+2)^3}\]

OpenStudy (anonymous):

a

OpenStudy (amistre64):

\[\sqrt[5]{x+8}=8\] \[(\sqrt[5]{x+8})^{1/5}=8^{1/5}\] \[x+8=8^{1/5}\] and continue

OpenStudy (amistre64):

pfft, ^5 not ^1/5

OpenStudy (amistre64):

for example: \[\sqrt[n]{b}=b^{1/n}\] if we ^n it we get \[(b^{1/n})^n=b^{n/n}=b\]

OpenStudy (amistre64):

so, x+8 = 8^5 and continue

OpenStudy (anonymous):

ok, so it's ok that the answer ends up being a crazy number? 32760?

OpenStudy (amistre64):

another way to prolly see it better is: \[\sqrt[5]{k}=8~~\implies~~k=8^5\]

OpenStudy (amistre64):

yes, big numbers are allowed to be solutions :)

OpenStudy (amistre64):

x = 8^5 - 8 is a find solution 3(5rt(8^5-8 +8)) + 3 = 27 3*5rt(8^5) + 3 = 27 3*8 + 3 = 27 is true

OpenStudy (anonymous):

thank you so much for not just giving me the answers and actually helping me understand how this works!

OpenStudy (amistre64):

youre welcome